OpenStudy (anonymous):

Describe the usefulness of the conjugate and its effect on other complex numbers

OpenStudy (anonymous):

\[\frac{ 2 }{ 4-i }* \frac{ 4+i }{ 4+i } = \frac{ 8+2i }{ 16+1 } = \frac{ 8+2i }{ 17 }\]

OpenStudy (anonymous):

It gets rid of the complex number on the bottom.

OpenStudy (anonymous):

If you are trying to find the magnitude of some value, it is much easier to solve if you have the complex part in the numerator.